OpenStudy (anonymous):

Factor completely: 35x^3-225x^2+250x Step by step please

OpenStudy (saifoo.khan):

go for commons.

OpenStudy (saifoo.khan):

5x(7x^2 -45x + 50) Now factor the brackets.

OpenStudy (saifoo.khan):

\[5x(x-5) (7 x-10) \]

OpenStudy (anonymous):

@saifoo.khan how did you factor inside the bracket?

OpenStudy (chaise):

\[35x^3-225x^2+250x\] \[=5x(7x^2-45x+50)\] \[=5x(x-5)(7x-10)\]

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=_BGAOjZzeCI

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=AMEau9OE6Bs

OpenStudy (anonymous):

Basically, what can you multiply to get the last number, and with these same numbers, add to get the middle number?

OpenStudy (anonymous):

Also if there's a coefficient in front of the x^2 then it's mostly guess and check. That's my way.

OpenStudy (anonymous):

@brandon376 if it's like you said how would -5 + -10 = -45? o_o

OpenStudy (chaise):

There is a method to solving them - factors of 7 and factor of 50 which cross multiply and then add to give -45. Difficult process made easy by the calculator.

OpenStudy (anonymous):

There's a 7 though

OpenStudy (anonymous):

@chaise can you please explain more

OpenStudy (anonymous):

x(35x^2-225x+250) =5x(7x^2-45x+50) =5x(7x^2-35x-10x+50) =5x(x-5)(7x-10)